OpenStudy (anonymous):

Find the exact values of each of the remaining trig functions of theta. sin(theta)=(-3/5) , 180*

OpenStudy (anonymous):

Quadrant III, x=(-4/5)

OpenStudy (anonymous):

cos = (-4/5) tan = (3/4) cot = (4/3) sec = (-5/4) csc = (-5/3)

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

Yes, those are all correct mikejano

OpenStudy (anonymous):

Thanks : )

OpenStudy (anonymous):

x=-4, y=-3, r=5, agree, all correct